Tohban Report 2015-06-17

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 10 Jun 2015
End Date: 17 Jun 2015
Tohban: Mitsuo Oka
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

Solar activity was higher than previous weeks with many C-class flares and 3 M-class flares. There are currently 7 ARs on the disk and 3 of them are expected to rotate off the disk within the next day or so. More C-class flares and isolated M-class flares are expected during the next few days.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    70     3     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    54     3     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0    22     0     0

There were RHESSI flares/GOES flares 409 / 73 over the time range 10-Jun-15 17-Jun-15

Memory Management

The SSR fill level reached a maximum of ~50% and a minimum of about ~26% during the week.

Spacecraft Status

The cold plate 1 and 2 are at about 127.2 K and 125.7 K respectively.

Data Gaps

There were data gaps (as well as noisy data) due to the replays of data.


GAP START TIME              GAP END TIME                   GAP (SEC)

2015-06-12T22:25:00.000 -- 2015-06-12T22:40:00.000       900.00000

2015-06-13T23:35:00.000 -- 2015-06-13T23:45:00.000       600.00000
2015-06-13T23:50:00.000 -- 2015-06-13T23:55:00.000       300.00000

2015-06-14T00:10:00.000 -- 2015-06-14T00:25:00.000       900.00000
2015-06-14T00:30:00.000 -- 2015-06-14T00:40:00.000       600.00000
2015-06-14T01:10:00.000 -- 2015-06-14T01:15:00.000       300.00000
2015-06-14T01:25:00.000 -- 2015-06-14T01:45:00.000       1200.0000
2015-06-14T02:00:00.000 -- 2015-06-14T02:10:00.000       600.00000
2015-06-14T05:40:00.000 -- 2015-06-14T05:55:00.000       900.00000

2015-06-15T15:10:00.000 -- 2015-06-15T16:20:00.000       4200.0000
2015-06-15T21:15:00.000 -- 2015-06-15T23:55:00.000       9600.0000

2015-06-16T02:35:00.000 -- 2015-06-16T02:45:00.000       600.00000
2015-06-16T07:35:00.000 -- 2015-06-17T00:00:00.000       59100.000

Detector issues

D8 Front fast-threshold from 0x22 to 0x30 at 2015-162-20:18:18

D9 Front fast-threshold from 0x40 to 0x50 at 2015-162-20:18:44

Other notes

Switched from normal_vigorous to active_vigorous at 2015-166-20:27:12 UTC (Jun 15, Monday). There has been stale TLE for about 4 days, leading to unstable tracking and many replays of data. It was considered that a change to "active_vigorous" would be a quick way to slow down the SSR filling up and allow Ops more time to replay the data.

There was a discussion on optimizing the threshold for shutter motions.

Spacecraft Management

Decimation Normal/Vigorous
Night time data (fronts) Taking night events
Night time data (rears) Taking night events
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Personal tools